8. The US Department of Agriculture reports that the mean cost of raising a child from birth to age 2 in a rural area is $10,460 with a standard deviation of $1,540. You believe the value of the mean cost is incorrect. So you want to test the claim by selecting a random sample of 900 children (age 2) and find that the mean cost is $10,345. At the 5% level of significance, is there enough evidence to conclude that the mean cost is different from $10,460?
Answer #1

8

Null and alternative hypotheses:

Ho : μ 10.460

Ha : μメ10.460

Since population standard deviation is given, we will use z test here.

Test statistic

10460-224 10345-10460 : = =- 1540/V900

p - Value = 2*P(z < -2.24) = 0.0251

Since this is less than 0.05, we reject Ho.

There is enough evidence to reject the claim that the mean cost of raising a child from birth to age 2 in a rural area is $10460.